OpenStudy (anonymous):

Factor the polynomial by grouping. 4a^3+8a^2+3a

OpenStudy (saifoo.khan):

a is the GCF.

OpenStudy (anonymous):

=a( 4a2 + 8a +3) =a( 4a2 + 6a+2a +3) =a(2a(2a+3)+(2a+3)) =a(2a+3)(2a+1)

OpenStudy (anonymous):

thanks